OpenStudy (anonymous):

I'm confused. Solve for x: log(3)(x)=3

OpenStudy (anonymous):

\[\log_{3}x=3 \]

OpenStudy (anonymous):

the base is 3

OpenStudy (anonymous):

x = log(base b) of y is equivalent to y = b^x

OpenStudy (anonymous):

\[Log_a(x) = b\] \[a^b = x\]

OpenStudy (anonymous):

so 9 =x?

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

x=1

OpenStudy (amistre64):

not 3*3; 3^3

OpenStudy (anonymous):

Thanks everyone!!!:)

OpenStudy (anonymous):

oh so 729

OpenStudy (anonymous):

No the answer is wrong... Sorry..

OpenStudy (amistre64):

3^3 = 3*3*3

OpenStudy (anonymous):

\[x = 3^3 = 3 \times 3 \times 3 = ???\]

OpenStudy (anonymous):

what the heck, lol I meant 27 idk why I put 729

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

the answer is x=1

OpenStudy (anonymous):

uhhhhhhhh, how in the world did you manage to get that answer?

OpenStudy (anonymous):

you said Solve for x log(3)(x)=3 3 x X = 3 BECAUSE X=1 SO 3X1=3 SO YOUR ANSWER IS = 3. SO TELL ME HOW YOU GET THE ANSWER =3 WHAT IS YOUR X

OpenStudy (anonymous):

YES THE CORRECT ANSWER IS X= 1

OpenStudy (anonymous):

Why are you yelling?

OpenStudy (amistre64):

your best bet is just to walk away and let them live in their own delusions ..... imo